Salem and vernon are 144 miles apart. A car leaves salem traveling towards vernon, and another car leaves vernon at the same time, traveling towards salem. The car leaving salem avergaes 10 miles per hour more than the other, and they meet after 1 hour and 36 minutes. What are the avergae speeds of the carfs?
Answer by mananth(16949)

car leaving vernon x mph
car leaving Salem = x+10 mph
..
they both leave towards each other ..
Add up speeds
x+x+10 = 2x +10 mph
..
time = 1 . 36 mins
= 1 3/5 hours
=8/5 hours
..
d= rt
144 = (2x+10)* 8/5
multiply by 5
5*144 = 8*(2x+10)
/8
5*144/8 = 2x+10
5*18 = 2x+10
90=2x+10
-10....... -10
90-10=2x
80=2x
/2
40 =x
x = 40 mph
other car 50 mph
...
m.ananth@hotmail.ca